OpenStudy (anonymous):

Set up and evaluate an integral to calculate the volume of the solid obtained by rotating the region between y = 5-x and y = 25-x^2 around the x-axis.

OpenStudy (anonymous):

so far i have \[\int\limits_{-4}^{5}(25-x^2)^2- (5-x)^2\]

OpenStudy (anonymous):

multiply it out to simplify it

OpenStudy (anonymous):

Expand[(25 - x^2)^2] - Expand[(5 - x)^2] = 600 + 10 x - 51 x^2 + x^4